# 2]Dept. of Physics and Astronomy and LENS, Florence Univ., via Sansone 1, I-50019 Sesto Fiorentino, Italy

### v1

@@ Abstract
Quantum machine learning is a new research field combining quantum information science and machine learning. Quantum computing technologies seem to be particularly well suited to solving problems in the health sector in an efficient way, because they may deal with large datasets more efficiently than classical AI.
Alzheimer’s disease is a neurodegenerative brain disorder that mostly affects elderly people, causing important cognitive impairments. It is the most common cause of dementia and it has an effect on memory, thought, learning abilities and movement control. This type of disease has no cure, consequently an early diagnosis is fundamental for reducing its impact. The analysis of handwriting can be effective for diagnosing, as many researches have conjectured. The DARWIN (Diagnosis AlzheimeR WIth haNdwriting) dataset contains handwriting samples from people affected by Alzheimer’s disease and a group of healthy people. Here we apply quantum AI to this use-case. In particular, we use this dataset to test kernel methods for classification task and compare their performances with the ones obtained via quantum machine learning methods. We find that quantum and classical algorithms achieve similar performances and in some cases quantum methods perform even better.
@@ 1 Introduction
Neurodegenerative diseases are incurable conditions caused by the progressive degeneration of nerve cells [1]. Alzheimer’s disease (AD) is one of the most common amongst them [2]. The predominant symptom in the early phase of AD is episodic memory impairment followed by progressive amnesia, a result of widespread brain damage. Since there is no cure for AD, it is critical to improve the approaches now used for diagnosis [3]. Handwriting is a task that involves both cognitive and motor functions to plan and properly execute the movements that are required with high coordination [4]. The analysis of handwriting and drawing dynamics is a valid non-invasive way for evaluating AD progression. For this purpose a protocol of 25 tasks was defined in Ref. [5]. Each handwriting/drawing task is described by 18 features and the data collected amongst a group of healthy people and patients made up the DARWIN (Diagnosis AlzheimeR WIth haNdwriting) dataset. In Ref. [6] it is described how this dataset can be exploited by machine learning models to discriminate between AD patients and healthy people.
@@
Machine learning (ML) is the field of artificial intelligence (AI) that studies the development of learning models to recognize, for example, patterns in large sets of data [7]. These models can learn how to attribute outputs to unseen inputs for classification or regression tasks (supervised learning), or find a structure in the inputs (unsupervised learning) for clustering or dimensionality reduction [8]. A lot of work has been done in the past years to improve the performance of ML algorithms and make them capable of handling efficiently big datasets. One more recent and promising solution could be to use quantum machine learning (QML) techniques. In QML inputs are encoded in quantum states and computation is done on a quantum computer [9, 10]. Although quantum processing units (QPUs) are still subjected to relatively large noise rates, it seems that applications in ML can lead to remarkable results. For example, in Ref. [11] it is shown how variational quantum models, namely models based on updating parameters in a quantum circuit through classical optimization, can exhibit a significant learning advantage in solving a regression task with input data from the fashion-MNIST dataset. A review of the contributions of QML in medical image analysis is given in Ref. [12]. A previous study [13] approaches via a so-called variational quantum classifier (VQC) the problem of classification defined by the same dataset we consider in our work.
@@
Here we focus in particular on quantum kernel methods, i.e. kernel methods in which the feature space is a space of quantum states. We use these algorithms to classify the samples and compare their performances with the ones achieved by classical algorithms. The paper is organized as follows. In Section 2 we describe how the DARWIN dataset was created and we define the classical and quantum kernel methods for classification. Then in Section 3 and 4 we report their performances, we discuss the results and draw our conclusions.
@@ 2.1 Dataset
The DARWIN dataset is composed of handwriting data collected according to the protocol defined in Ref. [5]. It includes 25 tasks that can be divided in the following categories: graphic tasks, like joining points and drawing geometrical figures; copy tasks, testing the ability of repeating more complex gestures to write letters, numbers and words; memory tasks, involving writing previously memorized words; and dictation tasks. The protocol was submitted to 174 participants: 89 AD patients and 85 healthy people, recruited so that the two groups would match in terms of age, education and gender. None of the participants was taking medications that influenced their cognitive abilities. The tasks were performed on paper sheets put on top of a tablet equipped with a pen that could both write in ink and sample coordinates of the tip and pressure exerted. The tablet was connected to a PC that recorded movements and displayed them in real-time. A software was used to extract 18 features for each task. Features and tasks are described in details in Appendix A. The dataset includes also a feature for identification of participants and another one that indicates whether the sample is associated to an AD patient (P) or an healthy person (H). This feature is used as label in the classification problem via a (classical or quantum) supervised learning model.
2.2 Support vector classification (SVC)
2.2.1 Classical SVC
Support vector machines (SVMs) are indeed supervised learning models for classification and regression [14]. Considering a binary classification problem in which the aim is to decide whether a set of samples belongs to one of two classes, each data point is viewed as a $d$-dimensional vector. The goal is to separate the points with a $(d-1)$-dimensional hyperplane such that the distance from the hyperplane to the nearest data points on each side, called support vectors, is maximized. A graphic representation of this is given in Fig. 2. In the original space this problem has often no solution, i.e. the sets of point are not linearly separable. A SVM uses the kernel trick, the technique of mapping inputs into high-dimensional feature spaces, to solve this problem [15]. The dot products of input pairs in the feature space are called kernels. Defining a map $\varphi:\mathcal{X}\rightarrow\mathcal{F}$ from the original input space to the feature space is equivalent to defining a kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{C}$ that maps couples of inputs into their dot products in the feature space.
@@ 2.2.1 Classical SVC
For our classification task we consider 3 different kernel functions [16]. The radial basis function (RBF) kernel maps a pair of inputs $(x,x^{\prime})$ into $k_{rbf}(x,x^{\prime})=\exp\left(-\gamma|x-x^{\prime}|^{2}\right).$ This is one of the most common choices amongst kernel functions and it is used by default in the standard version of SVC implemented in the Scikit library in Python [17]. The linear kernel function is a simple dot product, $k_{lin}(x,x^{\prime})=\langle x,x^{\prime}\rangle.$ Lastly, the sigmoid kernel is given by $k_{sig}(x,x^{\prime})=\tanh(\gamma\langle x,x^{\prime}\rangle+r)$.
2.2.2 Quantum SVC
@@
{quantikz}
[…]
In Fig. 3 we show the graphic representation of our 6-qubits circuit, composed by layers of $R_{x}$ and $R_{y}$ gates alternating with layers of controlled-z and controlled-x gates. On the right-hand side are symbols representing measurements applied to the qubits, converting the quantum states in a sequence of classical bits. The number of circuit parameters is 24.
[…]
Figure 3: Graphic representation of the 6-qubits PQC that maps a sample $s=(s[0],s[1],...,s[23])$ into a quantum state. The parameters in the rotation gates are given by a bandwidth factor $b=0.4$, chosen in order to maximize mean accuracy, times the components of $s$.
A similar structure with the same number of parameters is used to define also an 8-qubits and a 12-qubits circuit. The number of parameters must be equal to the number of features of the samples, so that each input feature corresponds to a parameter. Therefore the dataset needs to be processed through a principal component analysis (PCA) prior to applying the methods [25]. In order to reduce expressibility, the parameters in the rotation gates are multiplied by a bandwidth hyperparameter, chosen so as to maximize the accuracy reached by the classification. These circuits produce kernel matrices whose eigenvalue curve are plotted in Fig. 4. The curves are not flat but decay, hence the problem can be solved in this way.
@@ 2.3 Implementation
Our algorithms are implemented in Python, using the functionalities of Scikit and Qiskit libraries [26]. First we perform some preprocessing on the dataset. In particular we drop the identification feature, while keeping the feature healthy/AD’s patient (H/P) as a label. We standardize the samples, by subtracting the mean and scaling to unit variance. Then we perform a PCA with 24 components. Finally we standardize the set of samples derived from the PCA. In fact, from our analysis it emerges that this second normalization improves the performance of both classical and quantum methods. To reduce the bias introduced by selecting the samples for training and test, we select randomly 20 different splittings of the dataset in training and test set, keeping for the latter 21% of the samples, so that 137 samples are used for training and 37 for test. We use the same splittings for all methods.
After that we run classical SVCs using 3 different kernel functions: a radial basis function, a linear function and a sigmoid function. Subsequently we run the quantum SVCs for each of the 3 PQCs. Besides imposing the choice of the kernel function in the classical SVCs and the option of precomputing the kernels in the quantum SVCs, we keep the default setting for the other options of command ‘SVC’. For each run we compute the accuracy score, i.e. the number of test samples labelled correctly divided by the total number of test samples, and finally we calculate and compare the obtained average accuracies.
Table 1: Accuracy (in percentage) of classical SVCs and quantum SVCs for 20 different splittings of the dataset in training and test set. The highest accuracy in each run is reported in red.
Classical SVC Quantum SVC
Rbf Linear Sigmoid 6 qubits 8 qubits 12 qubits
81.08 86.49 86.49 86.49 83.78 89.19
83.78 81.08 86.49 83.78 81.08 81.08
75.68 67.57 83.78 78.38 75.68 75.68
94.59 89.19 81.08 91.89 91.89 91.89
83.78 81.08 81.08 81.08 86.49 83.78
89.19 83.78 91.89 94.59 89.19 91.89
91.89 78.38 89.19 91.89 94.59 91.89
91.89 83.78 89.19 89.19 91.89 89.19
89.19 89.19 86.49 83.78 91.89 89.19
94.59 97.30 91.89 89.19 94.59 94.59
86.49 86.49 86.49 89.19 83.78 86.49
94.59 91.89 86.49 91.89 94.59 91.89
86.49 83.78 78.38 81.08 86.49 86.49
78.38 83.78 83.78 81.08 86.49 89.19
86.49 75.68 75.68 81.08 81.08 81.08
86.49 89.19 89.19 89.19 94.59 89.19
86.49 86.49 81.08 81.08 89.19 86.49
83.78 86.49 81.08 86.49 91.89 86.49
86.49 83.78 83.78 83.78 89.19 86.49
83.78 83.78 89.19 83.78 91.89 91.89
@@ 3 Results and Discussion
In Table 1 and in Fig. 5 we show the accuracy reached by the methods for each of the 20 splittings of the dataset in training and test set. Amongst classical SVCs, the one defined by the RBF kernel achieves average accuracy of 86.76%. Linear kernel and sigmoid kernel obtain 84.46% and 85.14% respectively. The quantum SVC defined by the 6-qubits PQC gets to 85.95%, the one defined by the 8-qubits PQC reaches 88.51% and the one defined by the 12-qubits circuit achieves average accuracy of 87.70%. One of the classical methods performs better than any other quantum method for 5 splittings. In 3 other splittings the same best precision is achieved by a classical and a quantum method. In the other 12 cases the quantum methods perform better. In particular, the 6-qubits method performs best in 2 splittings, the 8-qubits method in 7 splittings, the 12-qubits method achieves the best precision in 2 cases, while it obtains the same best accuracy of the 8-qubits method in 1 other case.
Figure 5: Accuracies (in percentage) vs. splittings of the dataset in classical and quantum SVC models.
@@ 3.1 Data subsampling
Now we repeat the analysis above when restricted to only one category of features, e.g. the features derived only from graphic tasks or only copy tasks or only memory and dictation tasks. We compare the performance of the classical and quantum SVCs that have achieved the highest average accuracies above, i.e. the classical SVC defined by the RBF kernel and the 8-qubits quantum SVC.
Table 2: Accuracies (in percentage) of RBF classical SVC and 8-qubits SVC achieved combining with the majority vote decision rule the predictions made considering all tasks individually (left) and considering only the best 5 tasks (right). The highest accuracy in each run is reported in red.
All tasks Best 5 tasks
Classical Quantum Classical Quantum
78.38 81.08 89.19 91.89
81.08 78.38 86.49 86.49
72.97 72.97 86.49 83.78
78.38 81.08 91.89 91.89
81.08 81.08 91.89 91.89
78.38 81.08 81.08 83.78
89.19 89.19 86.49 86.49
72.97 75.68 81.08 83.78
83.78 86.49 91.89 91.89
91.89 94.59 94.59 94.59
89.19 83.78 89.19 89.19
70.27 72.97 89.19 89.19
75.68 72.97 81.08 83.78
78.38 78.38 81.08 81.08
81.08 81.08 78.38 78.38
78.38 83.78 83.78 83.78
89.19 91.89 89.19 86.49
86.49 83.78 94.59 89.19
75.68 78.38 89.19 89.19
75.68 81.08 86.49 86.49
We use the same 20 splittings of the dataset in training and test set we have used above. First we take into account the features derived from the tasks in the category ‘graphic’. RBF and 8-qubits SVC achieve average accuracies of 79.19% and 79.86% respectively. After that we consider only the features derived from the tasks in the category ‘copy’. Both classical and quantum SVC achieve average accuracies of 83.11%. Finally, we focus on the features derived from the tasks in the category ‘memory and dictation’. Classical and quantum SVC achieve mean accuracies of 79.46% and 78.38%.
Subsequently we consider the features obtained from each of the 25 task individually. For each task we run RBF SVC, store the predictions made by the 25 models and use the majority vote decision rule [27] to select the outputs; namely for each sample we select the output predicted by most of the models. Likewise we run for each task a quantum SVC. Since each task is described by 18 features, we need to define a new PQC with 18 parameters. For this purpose we choose a 9-qubits circuit. No PCA is necessary in this case, since the number of parameters coincides with the number of features. We perform 20 runs with the same splittings in training and test set we have used above. The accuracies reached at each run are shown in Table 2. Classical and quantum SVC obtain mean accuracies of 80.41% and 81.49% respectively. Afterwards we select the best 5 tasks, namely the tasks that have achieved singularly the best accuracies in the previous runs. For both classical and quantum SVC the best tasks, although in a different order, are 23, 19, 17, 24 and 8. The tasks are described in Appendix A. As before we use the majority vote decision rule to select the predictions amongst those made by the 5 models associated with the tasks. In this case both classical and quantum SVC obtain average accuracies of 87.16%.

### v2

@@ Abstract
Quantum machine learning is a new research field combining quantum information science and machine learning. Quantum computing technologies appear to be particularly well-suited for addressing problems in the health sector efficiently. They have the potential to handle large datasets more effectively than classical models and offer greater transparency and interpretability for clinicians.
Alzheimer’s disease is a neurodegenerative brain disorder that mostly affects elderly people, causing important cognitive impairments. It is the most common cause of dementia and it has an effect on memory, thought, learning abilities and movement control. This type of disease has no cure, consequently an early diagnosis is fundamental for reducing its impact. The analysis of handwriting can be effective for diagnosing, as many researches have conjectured. The DARWIN (Diagnosis AlzheimeR WIth haNdwriting) dataset contains handwriting samples from people affected by Alzheimer’s disease and a group of healthy people. Here we apply quantum AI to this use-case. In particular, we use this dataset to test classical methods for classification and compare their performances with the ones obtained via quantum machine learning methods. We find that quantum methods generally perform better than classical methods.
@@ 1 Introduction
Neurodegenerative diseases are incurable conditions caused by the progressive degeneration of nerve cells [1]. Alzheimer’s disease (AD) is one of the most common amongst them [2]. The predominant symptom in the early phase of AD is episodic memory impairment followed by progressive amnesia, a result of widespread brain damage. Since there is no cure for AD, it is critical to improve the approaches now used for diagnosis [3]. Early diagnosis can make patients eligible for clinical trials, which are essential for the development of new treatments. Participation in these trials can provide access to cutting-edge therapies that are not yet widely available [4] and it can help maintain cognitive function for a longer period, allowing individuals to live more independently and enjoy a better quality of life [5].
@@
Many techniques have been used for this purpose. Neuroimaging techniques are commonly used to detect brain atrophy and other structural changes associated with AD; however their limitations include restricted availability and the inability to detect early-stage diseases [6]. Cognitive assessments, on the other hand, despite potentially being influenced by the patient’s education level, measure abilities such as memory, attention, language, and problem-solving, while motor tests examine coordination, balance, and gait. This comprehensive approach aids in fully understanding the extent of the disease’s impact [7]. Handwriting is a task that involves both cognitive and motor functions to plan and properly execute the movements that are required with high coordination [8]. The analysis of handwriting and drawing dynamics is a valid non-invasive way for evaluating AD progression. For this purpose a protocol of 25 tasks was defined in Ref. [9]. Each handwriting/drawing task is described by 18 features and the data collected amongst a group of healthy people and patients made up the DARWIN (Diagnosis AlzheimeR WIth haNdwriting) dataset. In Ref. [10] it is described how this dataset can be exploited by machine learning models to discriminate between AD patients and healthy people.
@@
Machine learning (ML) is a subfield of artificial intelligence (AI) that focuses on the development of algorithms and statistical models that enable computers to learn and make predictions or decisions based on data [11]. These models can identify patterns, trends, and relationships within large datasets, allowing for tasks such as classification, regression, clustering, and anomaly detection [12]. One of the main obstacles to the usage of AI in the healthcare field, and in particular for the diagnosis of diseases, is the “black box” nature of many AI models. Medical practitioners often find it challenging to trust AI predictions because they cannot easily understand how these models arrive at their conclusions [13]. Quantum AI can improve the transparency and interpretability of AI models by leveraging the principles of quantum mechanics and providing a clear visual representation of models [14], allowing clinicians to understand the decision-making process better. Additionally, it can handle large volumes of data more efficiently than classical AI, making it a powerful tool for data-intensive tasks [15]. In quantum machine learning (QML) inputs are encoded in quantum states and computation is done on a quantum computer [16, 17]. Although quantum processing units (QPUs) are still subjected to relatively large noise rates, it seems that applications in ML can lead to remarkable results. For example, in Ref. [18] it is shown how variational quantum models, namely models based on updating parameters in a quantum circuit through classical optimization, can exhibit a significant learning advantage in solving a regression task with input data from the fashion-MNIST dataset. A review of the contributions of QML in medical image analysis is given in Ref. [19]. A previous study [20] approaches via a variational quantum classifier (VQC) the problem of classification defined by the same dataset we consider in our work.
@@
Here we focus in particular on quantum kernel methods, i.e. kernel methods in which the feature space is a space of quantum states. We utilize these algorithms to classify the samples and evaluate their performance against a classical kernel algorithm and two additional classical methods. The paper is organized as follows. In Section 2 we describe how the DARWIN dataset was created and we define the classical and quantum methods for classification. Then in Section 3 and 4 we report their performances, we discuss the results and draw our conclusions.
@@ 2.1 Dataset
The DARWIN dataset is composed of handwriting data collected according to the protocol defined in Ref. [9]. This protocol includes 25 tasks that can be divided in the following categories: graphic tasks, like joining points and drawing geometrical figures; copy tasks, testing the ability of repeating more complex gestures to write letters, numbers and words; memory tasks, involving writing previously memorized words; and dictation tasks. The protocol was submitted to 174 participants: 89 AD patients and 85 healthy people, recruited so that the two groups would match in terms of age, education and gender. None of the participants was taking medications that influenced their cognitive abilities. The tasks were performed on paper sheets put on top of a tablet equipped with a pen that could both write in ink and sample coordinates of the tip and pressure exerted. The tablet was connected to a PC that recorded movements and displayed them in real-time. A software was used to extract 18 features for each task. Features and tasks are described in details in Appendix A. The dataset includes also a feature for identification of participants and another one that indicates whether the sample is associated to an AD patient (P) or an healthy person (H). This feature is used as label in the classification problem via a (classical or quantum) supervised learning model.
2.2 Classical SVC
Support vector machines (SVMs) are supervised learning models for classification and regression [21]. Considering a binary classification problem in which the aim is to decide whether a set of samples belongs to one of two classes, each data point is viewed as a $d$-dimensional vector. The goal is to separate the points with a $(d-1)$-dimensional hyperplane such that the distance from the hyperplane to the nearest data points on each side, called support vectors, is maximized. A graphic representation of this is given in Fig. 2. In the original space this problem has often no solution, i.e. the sets of point are not linearly separable. A SVM uses the kernel trick, the technique of mapping inputs into high-dimensional feature spaces, to solve this problem [22]. The dot products of input pairs in the feature space are called kernels. Defining a map $\varphi:\mathcal{X}\rightarrow\mathcal{F}$ from the original input space to the feature space is equivalent to defining a kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{C}$ that maps couples of inputs into their dot products in the feature space.
@@ 2.2 Classical SVC
For our classification task we consider different kernel functions [23]. The radial basis function (RBF) kernel maps a pair of inputs $(x,x^{\prime})$ into $k_{rbf}(x,x^{\prime})=\exp\left(-\gamma|x-x^{\prime}|^{2}\right).$ This is one of the most common choices amongst kernel functions and it is used by default in the standard version of SVC implemented in the Scikit library in Python [24]. The linear kernel function is a simple dot product, $k_{lin}(x,x^{\prime})=\langle x,x^{\prime}\rangle.$ The polynomial function is $k_{pol}(x,x^{\prime})=(\gamma\langle x,x^{\prime}\rangle+r)^{d}.$ Lastly, the sigmoid kernel is given by $k_{sig}(x,x^{\prime})=\tanh(\gamma\langle x,x^{\prime}\rangle+r)$.
2.3 k-Nearest Neighbors
The k-Nearest Neighbors (kNN) model is a method used both for classification and regression [25]. With this technique, a label is assigned to a data point based on the most common label among its $k$ nearest neighbors. In particular, the probability that the unknown sample $x_{0}$ belongs to class $j$ is given by
$P(Y=j\mid X=x_{0})=\frac{1}{k}\sum_{i\in N_{0}}I(y_{i}=j),$
where $k$ is the number of nearest neighbors, $N_{0}$ is the set of the $k$ nearest neighbors to $x_{0}$ and $I(y_{i}=j)$ is an indicator function that equals 1 if the $i$-th neighbor $y_{i}$ has the label $j$, and 0 otherwise.
2.4 Decision Tree
A Decision Tree (DT) is a decision support tool that visually represents decisions and their possible outcomes in a hierarchical, tree-like structure [26]. This structure is composed of:
  • nodes, which are decision points where a specific feature is tested;
  • branches, representing the possible outcomes of the test conducted at each node;
  • leaf nodes where a final decision is made and a label is assigned.
The tree is constructed by recursively splitting the data based on the attribute that provides the best separation according to a chosen criterion (e.g., Gini impurity or information gain). This process continues until the stopping criteria are met, such as a maximum tree depth or a minimum number of samples per leaf node.
2.5 Quantum SVC
@@
In Fig. 3 we show the graphic representation of a 6-qubits circuit, composed by layers of $R_{x}$ and $R_{y}$ gates alternating with layers of controlled-z and controlled-x gates. On the right-hand side are symbols representing measurements applied to the qubits, converting the quantum states in a sequence of classical bits. The number of circuit parameters is 24.
Figure 3: Graphic representation of a 6-qubits PQC that maps a sample $s=(s[0],s[1],...,s[23])$ into a quantum state. The parameters in the rotation gates are given by a bandwidth factor $b=0.4$, chosen in order to maximize the accuracy, times the components of $s$.
A similar structure with the same number of parameters is used to define also an 8-qubits and a 12-qubits circuit. The number of parameters must be equal to the number of features of the samples, so that each input feature corresponds to a parameter. Therefore the dataset needs to be processed through a principal component analysis (PCA) prior to applying the methods [34]. In order to reduce expressibility, the parameters in the rotation gates are multiplied by a tunable bandwidth hyperparameter. These circuits produce kernel matrices whose eigenvalue curve are plotted in Fig. 4. Since the curves decay, the problem can be addressed using this approach.
@@ 2.6 Implementation
Our algorithms are implemented in Python, using the functionalities of Scikit and Qiskit libraries [35]. First we perform some preprocessing on the dataset. In particular we drop the identification feature, while keeping the feature healthy/AD’s patient (H/P) as a label. We standardize the samples, by subtracting the mean and scaling to unit variance. Then we perform a PCA with 24 components. Finally we standardize the set of samples derived from the PCA. In fact, from our analysis it emerges that normalizing twice improves the performance of both classical and quantum methods. The dataset is divided for cross-validation using ShuffleSplit, performing 20 splits: 60% of the data is used for training, 20% for validation and 20% as test set. For each method, a parameter grid is defined for hyperparameter tuning: in this phase the model is trained on the training set and evaluated on the validation set. After that, the model with the best hyperparameters is trained on the union of training and validation set and is evaluated on the test set. The mean and standard deviation of the accuracies across the splits are calculated.
For classical SVC, the parameter grid includes different options for 4 hyperparameters:
  • ‘kernel’, that defines the kernel function;
  • ‘C’, used to balance error minimization and model complexity;
  • ‘gamma’, which determines the influence range of training examples;
  • ‘tol’, which sets the precision for stopping criteria.
Two more classical methods for classifications are used for comparison, kNN and DT. For kNN the parameter grid includes 3 hyperparametrs: number of neighbors, metric and type of weights. For DT it includes 5 hyperparameters: separation criterion, splitter, maximum depth, minimum samples for split and minimum samples for leaf.
For quantum SVCs, we use fixed ansatze with 6, 8, and 12 qubits. To simplify the computations, the only hyperparameter tuned during the validation phase is the bandwidth parameter. Therefore, except for selecting the option of running the model with precomputed kernels, the other parameters of SVC() are left by default: $\mathrm{C}=1.0$, ‘gamma’=‘scale’, i.e. it scales with the number of samples, and ‘tol’=0.001. The quantum methods are simulated locally using Aer Simulator.
@@ 3 Results and Discussion
In Table 1 we show the mean accuracy and standard deviation of classical and quantum methods. Amongst classical methods, SVC has the highest mean accuracy, but the standard deviation is relatively high, suggesting that its performance can vary significantly across different splits; kNN and DT are less effective in terms of average performance, although the latter offers more consistent results. The 6-qubits and 8-qubits SVCs perform comparably to the classical SVC, with slightly lower accuracy but a smaller standard deviation, indicating more consistent performance. The quantum SVC with 12 qubits outperforms all other models, both classical and quantum, with the highest accuracy and best consistency. This result suggests that increasing the number of qubits can enhance the performance of quantum classifiers.
Table 1: Mean accuracy and standard deviation (expressed in percentage) of classical SVC, k-Nearest Neighbors, decision tree and quantum SVCs defined by 6-qubits, 8-qubits and 12-qubits PQCs.
Classical Quantum
SVC kNN DT 6q 8q 12q
$85.28\pm 7.46$ $69.57\pm 8.20$ $73.57\pm 5.26$ $83.57\pm 4.59$ $83.14\pm 4.51$ $88.29\pm 4.69$
@@ 3.1 Data subsampling
Now we compare the performance of classical and quantum methods when restricted to subsets of features. For simplicity, we consider only the classical SVC and 12-qubits quantum SVC. To properly evaluate the performance of the methods on feature subsets, we use for both methods the hyperparameters that have been most frequently chosen in the previous analysis. First we use one category of features at a time, e.g. the features derived only from graphic tasks, then only copy tasks, and finally only memory and dictation tasks. The dataset is divided for cross-validation using ShuffleSplit, performing 20 splits: 80% of the data is used for training and 20% for testing. In Table 2 we show the mean accuracy and standard deviation of classical and quantum methods.
Table 2: Mean accuracy and standard deviation (expressed in percentage) of classical SVC and 12-qubits SVC models applied to the subset of features obtained through copy, graphic or memory and dictation tasks.
Classical 12q
Copy $85.57\pm 8.00$ $85.71\pm 7.06$
Graphic $78.14\pm 4.97$ $81.29\pm 5.68$
Memory $79.28\pm 6.19$ $78.57\pm 4.99$
Subsequently we consider the features obtained from each of the 25 tasks individually. For each task we run classical SVC, store the predictions made by the 25 models and use the majority vote decision rule [36] to select the outputs; namely for each sample we select the output predicted by most of the models. Since each task is described by 18 features, we need to define a new PQC with 18 parameters. For this purpose we choose a 9-qubits circuit. After considering all 25 tasks, we select the 5 tasks that achieve singularly the best accuracies in the previous runs. For classical SVC the tasks are 21, 17, 16, 7, 23, while for quantum SVC they are 21, 17, 24, 14, 23. As before we use the majority vote decision rule to select the predictions amongst those made by the 5 models associated with the tasks. In Table 3 we show the mean accuracy and standard deviation of classical and quantum methods over 20 splittings.
Table 3: Mean accuracy and standard deviation (expressed in percentage) of classical SVC and 9-qubits SVC models achieved combining with the majority vote decision rule the predictions made considering all tasks individually and considering only the best 5 tasks.
Classical 9q
All 25 $85.71\pm 3.59$ $86.00\pm 4.30$
Best 5 $80.28\pm 4.23$ $81.35\pm 5.19$
Across all data subsampling scenarios, the quantum algorithm consistently demonstrates improved or comparable performance, underscoring its robustness and potential benefits in managing diverse data types.